# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

### v2

@@ III.1 Comparisons with previous results
Tab.(4) depict additional energy results for $n$=0.875 and $U$=8 under PBC. All results in the table are obtained by direct optimizations. For $t^{\prime}$=0 on the $8\times 16$ lattice, the vertical stripe order is obtained with the periods same as those on the $16\times 16$ lattice. Specifically, we investigate the case for non-zero $t^{\prime}$, which is promising to have superconductivity Hubbard_SC1 . For $t^{\prime}=-0.2$, we find that NNN backflow achieves slightly lower energy than that by NN backlfow, on the $8\times 16$ lattice. Thus, NNN backflow terms are considered for non-zero $t^{\prime}$.
@@
$8\times 24$0.9394 -0.9421
$8\times 32$0.9401 -0.9426
Table 5: Energy results obtained by direct optimizations for the filling $n$=0.8 and $n$=0.9375, with $t^{\prime}$=0 under PBC. On the $10\times 10$ lattice, all-site backflow has moderate improvement comparing to NN backflow for $n$=0.8 and $U$=8.
Figure 6: Charge and spin structure factors on the $10\times 20$ lattice for $n$=0.8 (a)(b) and on the $8\times 32$ lattice for $n$=0.9375 (c)(d), under PBC. (a)(b)For $n$=0.8, the spin structure factor peaks at $54\pi/$ 5 when $U$=8, however there are no obvious peaks when $U$=4. For both $U$=4,8, there are no obvious peaks in the charge structure factor. (c)(d)For $n$=0.9375, the spin structure factor peaks at $1615\pi/$ 16 when $U$=4, meanwhile there are no obvious peaks in the charge structure factor.
Fig.(5) depicts the ground state pattern for $t^{\prime}$=-0.2 on the $8\times 16$ lattice, with definitions of average spin and density per site from Fig.(2). From the figure, an obvious width-4 stripe is achieved, which yields to the recent AFQMC results on width-8 lattices Hubbard_SC1 . From results on larger lattices such as $12\times 12$ and $16\times 16$, there are no obvious patterns from $s_{i}^{z}$ or $n_{i}$. In Fig.(5), we compare spin structure factors between $t^{\prime}$=0 and -0.2. To reval detailed differences, we increase the number of k-points from $17^{2}$ to $65^{2}$. Thus, there are totally 33 k-points in the x-axis. From the figure, for $t^{\prime}$=0 peaks are sharply distrubuted and centered on $87\pi/$ 8. However, the peak for $t^{\prime}$=-0.2 is wider comparing to that for $t^{\prime}$=0. The wider peak reveals competitions between stripe orders of different periods in the ground state for $t^{\prime}$=-0.2, thus direct optimizations are easier to obtain local optimizations on larger lattices.
III.3 n=0.8 and n=0.9375
Energy results for $n$=0.8 and $n$=0.9385 are presented on Tab.(5). As depicted by cases of $n$=0.8, $U$=8 on the $10\times 10$ lattice, the improvement of all-site backflow is moderate comparing to NN backflow. Thus, the majority improvement of the state representation is achieved by NN backflow terms.
Fig.(6) depicts charge and spin structure factors for $n$=0.8 on the $10\times 20$ lattice and for $n$=0.9375 on the $8\times 32$ lattice. Based on the phase diagram via AFQMC AFQMC_phase_diagram , the critical interaction strength $U_{c}$ between un-ordered and ordered ground states increase with the doping level $h$. Here, for $n$=0.8, the spin structure factor peaks at $54\pi/$ 5 for $U$=8, meanwhile there are no obvious peaks for $U$=4. Comparing charge structure factors between $U$=4 and 8, further increasing $U$ will lead to the obvious peak in the charge structure factor. For $n$=0.9375, the spin structure factor peaks at $1615\pi/$ 16. Results of $n$=0.8 and 0.9375 obtained by direct optimizations yield to the phase diagram via AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
@@ IV Discussions and conclusions
Figure 7: Density-density correlations for $n$=0.875, $U$=8 on the $8\times 32$ lattice under PBC. Blue dots depict the $p$=0 wave-function and yellow dots depict the $p$=1 wave-function. For both $p$=0 and $p$=1, results are obtained based on 46080 MC samples, with the interval between two samples twice as the lattice size.
@@
In this work, we benchmarked the Tensor-Backflow method along with a Lanczos optimization for various fillings $n$, interaction strengths $U$, NNN strengths $t^{\prime}$, lattice sizes and boundary conditions. We investigate ground state properties through charge and spin structure factors, and results consistent with other state-of-the-art methods such as fPEPS for $n$=0.875 as well as AFQMC for $n$=0.8 and 0.9375 are obtained.
@@
From our results, although a Lanczos optimization significantly improves the energy, the property of the state is determined by the $p$=0 wave-function. Fig.(7) depicts density-density correlations from $p$=0 and $p$=1 wave-functions for $n$=0.875 and $U=8$ on the $8\times 32$ lattice under PBC. From the figure, the CDW with the period of 8 is already in the $p$=0 wave-function and there are no significant differences for $p$=1 results. Thus, we evaluate pair-pair correlations with definitions from literature pair_pair_correlation based on $p$=0 wave-functions for the case of $n$=0.875 and $U$=8 on the $16\times 16$ lattice. For either $t^{\prime}$=0 and -0.2, pair-pair correlations $P(d)$ are in the magnitude of $\mathcal{O}(10^{-5})$ when $d_{x}$=8 or $d_{y}$=8. Because energies with different orders are competitive when $n$=0.875, obtaining local optimizations are common under direct optimizations. Therefore, more careful and intensive investigations are necessary for $n$=0.875.
@@
Based on results in this work, the Tensor-Backflow is an efficient state representation for the latticed Fermi-Hubbard model. Furthermore, there are several ways to improve the energy precision. For example, considering more backflow terms rather than NN backflow to improve energy precision for the Hubbard model with NN hoppings, comparing energies under various initializations, applying more advanced optimization methods such as the stochastic reconfiguration, etc. Currently, the wave-function is not symmetry enforced, thus translational invariance is not preserved during optimizations. For PBC, the energy precision can be further improved by enforcing translational invariance on the wave-function. The Tensor-Backflow has no restrictions on lattice shapes and boundary conditions, thus it is potentially an efficient and universal method for solving latticed fermion systems.

### v3

@@ III.2 Comparisons with previous results
Tab.(5) depicts additional energy results for nn=0.875 and UU=8 under PBC for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. All results in the table are obtained by direct optimizations. For $t^{\prime}$=0 on the $8\times 16$ lattice, the vertical stripe order is obtained with same periods as those on the $16\times 16$ lattice.
In particular, energies for $t^{\prime}$=-0.2 are depicted in Tab.(5). The $t^{\prime}$=-0.2 is important for understanding mechanisms of superconductivity Hubbard_SC1 . From energy comparisons in Tab.(5), we find that NNN backflow (QQ=9) achieves slightly lower energy than NN backlfow (QQ=5) for the case of $t^{\prime}$=-0.2 on the $8\times 16$ lattice. Because of the non-zero NNN hopping strengths, NNN backflow are considered for all cases of $t^{\prime}$=-0.2.
Figure 5: Patterns of ground states for nn=0.875, UU=8 and $t^{\prime}$=-0.2 under PBC on the $8\times 16$ lattice, from the pp=1 wave-function with NNN backflow. Left-sided (right-sided) figure denotes the average spin (density) per site, within definitions from Fig.(3).
Figure 6: Comparisons of spin structure factors between $t^{\prime}$=0 and -0.2 for nn=0.875 and UU=8 under PBC. All results are from pp=1 wave-functions, with NN backflow for $t^{\prime}$=0 and NNN backflow for $t^{\prime}$=-0.2. Number of k-points are increased from $17^{2}$ to $65^{2}$ when calculating spin structure factors. The horizontal axis is $k_{x}$ for $t^{\prime}$=-0.2 and $k_{y}$ for $t^{\prime}$=0, meanwhile vertical axis is $S(k_{x},k_{y}=\pi)$ for $t^{\prime}$=-0.2 and $S(k_{x}=\pi,k_{y})$ for $t^{\prime}$=0.
Because cases of $t^{\prime}$=-0.2 are more challenging than those of $t^{\prime}$=0, optimizations are easily to get trapped on local minima, especially on large lattices. We show that choosing a proper initial UHF is very helpful to avoid the local minima for the final converged energy Hubbard_NN4 . In Tab.6, we compare energies converged by Tensor-Backflow with initial UHFs from different strengths of $U$ on the $12\times 12$ lattice. For cases of UU=8, $U_{\mathrm{UHF}}$=4 leads to the best converged energy of Tensor-Backflow. Ground state patterns on the $12\times 12$ lattice for various $U_{\mathrm{UHF}}$ are depicted in Appendix.D.
On the $8\times 16$ lattice, the ground state pattern for $t^{\prime}$=-0.2 is depicted in Fig.(5), with definitions of average spin and density per site from Fig.(3). For the case on the $8\times 16$ lattice, the initial UHF is insufficiently optimized under $U_{\mathrm{UHF}}=U$. From Fig.(5), a width-4 stripe is achieved, yielding to the result on the same lattice from NQS Hubbard_NN3 . In the optimization of the NQS, energies were compared between vertical and horizontal stripes and the horizontal stripe achieved lower energy Hubbard_NN3 . In this work, the horizontal stripe is achieved naturally after optimizations without pinning-fields.
In Fig.(6), we compare spin structure factors between $t^{\prime}$=0 and -0.2. To reval detailed differences, we increase the number of k-points from $17^{2}$ to $65^{2}$. Thus, there are totally 33 k-points in the x-axis. From the figure, for $t^{\prime}$=0 peaks are sharply distributed and centered on $7\pi/8$. However, the peak for $t^{\prime}$=-0.2 is wider comparing to that for $t^{\prime}$=0. The wider peak reveals competitions between stripe orders of different periods in the ground state for $t^{\prime}$=-0.2, thus the optimization is easily to get stuck on a local minima. The high optimization difficulty of $t^{\prime}$=-0.2 is depicted by Tab.(6) on the $12\times 12$ lattice. Therefore, comparisons of various $U_{\mathrm{UHF}}$ are necessary for $t^{\prime}$=-0.2 on the $16\times 16$ lattice in near future.
III.4 n=0.8 and n=0.9375
@@
$8\times 24$ -0.9394 -0.9421
$8\times 32$ -0.9401 -0.9426
Table 7: Energy results for the filling nn=0.8 and nn=0.9375, with $t^{\prime}$=0 under PBC. For all cases in the table, initial UHFs are insufficiently optimized with $U_{\mathrm{UHF}}=U$. On the $10\times 10$ lattice, all-site backflow has moderate improvement comparing to NN backflow for nn=0.8 and UU=8.
Figure 7: Charge and spin structure factors on the $10\times 20$ lattice for nn=0.8 and $t^{\prime}$=0. All results are from pp=1 wave-functions with NN backflow in wave-functions. (a)(b) and on the $8\times 32$ lattice for nn=0.9375 (c)(d), under PBC. (a)(b)For nn=0.8, the spin structure factor peaks at $4\pi/5$ when UU=8, however there are no obvious peaks when UU=4. For both UU=4,8, there are no obvious peaks in the charge structure factor. (c)(d)For nn=0.9375, the spin structure factor peaks at $15\pi/16$ when UU=4, meanwhile there are no obvious peaks in the charge structure factor.
For other fillings such nn=0.8 and nn=0.9375, energy results are presented on Tab.(7). As depicted by cases of nn=0.8, UU=8 on the $10\times 10$ lattice, the improvement of all-site backflow is moderate comparing to NN backflow. Thus, for a Hamiltonian with NN hoppings, the major state representation is achieved by the NN backflow.
Fig.(7) depicts charge and spin structure factors for nn=0.8 on the $10\times 20$ lattice and for nn=0.9375 on the $8\times 32$ lattice. Based on the phase diagram via AFQMC AFQMC_phase_diagram , the critical interaction strength $U_{c}$ between un-ordered and ordered ground states increase with the doping level $h$. Here, for nn=0.8, the spin structure factor peaks at $4\pi/5$ for UU=8, meanwhile there are no obvious peaks for UU=4. Comparing charge structure factors between UU=4 and UU=8, there is the tendency of the obvious peak in the charge structure factor when further increasing $U$. For nn=0.9375, the spin structure factor peaks at $15\pi/16$, meanwhile there are no obvious peaks in the charge structure factor. Results of nn=0.8 and nn=0.9375 in Fig.(7) yield to the phase diagram via AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
@@ IV Discussions and conclusions
In this work, we benchmarked the Tensor-Backflow method along with a Lanczos optimization for various fillings $n$, interaction strengths $U$, NNN hopping strengths $t^{\prime}$, lattice sizes and boundary conditions. Comparing to current state-of-the-art methods such as fPEPS Hubbard_fPEPS , NQS Hubbard_NN3 ; Hubbard_NN4 , achieved energies are competitive or even lower. In addition to energy results, we investigate ground state properties through patterns of $s_{i}^{z}$ and $n_{i}$ as well as charge and spin structure factors. When $t^{\prime}$=0, obtained state properties are consistent with other state-of-the-art methods such as fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 for nn=0.875 as well as AFQMC for nn=0.8 AFQMC_phase_diagram and nn=0.9375 AFQMC_n0.9375 . For $t^{\prime}$=-0.2 on the $8\times 16$ lattice, the width-4 horizontal stripe is achieved, matching the result from NQS Hubbard_NN3 . On the $12\times 12$ lattice, achieved energy is lower than the state-of-the-art NQS result Hubbard_NN3 , meanwhile a width-4 stripe is obtained based on $n_{i}$.
Figure 8: Density-density correlations for nn=0.875, UU=8 and $t^{\prime}$=0 on the $8\times 32$ lattice under PBC. With NN backflow in the wave-function. Blue dots depict the pp=0 wave-function and yellow dots depict the pp=1 wave-function. For both pp=0 and pp=1, results are obtained based on 46080 MC samples, with the interval between two samples twice as the lattice size.
@@
From our results, although a Lanczos optimization significantly improves the energy, the property of the state is determined by the pp=0 wave-function. Namely, a Lanczos optimization can’t override the property of the original wave-function, as patterns of $s_{i}^{z}$ and $n_{i}$ are similar between pp=0 and pp=1 wave-functions. For example, Fig.(8) depicts density-density correlations from pp=0 and pp=1 wave-functions for nn=0.875 and $U=8$ on the $8\times 32$ lattice under PBC. From the figure, the CDW with the period of 8 is already in the pp=0 wave-function and there are no significant differences for pp=1 results. Therefore, differences between correlations from pp=0 and pp=1 wave-functions are negligible.
Pair correlations determine the existence of superconductivitiy tJ_fPEPS ; pair_pair_correlation_1 ; pair_pair_correlation_2 . In this work, based on comparisons between pp=0 and pp=1 wave-functions, to avoid high calculation complexity, we evaluate $d$-wave pair correlations based on pp=0 wave-functions, with pair correlations defined in the literature tJ_fPEPS . For the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, when $t^{\prime}$=0 and $t^{\prime}$=-0.2, pair correlations are in the magnitude of $\mathcal{O}(10^{-5})$ when the distance between two pairs is $L/2$, where $L$ is the side length of the square lattice. Because optimizations are easily to get trapped on local minima for cases of $t^{\prime}$=-0.2, more careful and intensive investigations are necessary in the future work.
@@
Based on results in this work, the Tensor-Backflow is an efficient state representation for the lattice Fermi-Hubbard model. Furthermore, there are several ways to improve the energy precision. For example, considering more backflow terms comparing to hopping terms in the Hamiltonian, comparing energies under various initializations, comparing ground state patterns with and without applying pinning-fields, applying more advanced optimization methods such as the stochastic reconfiguration, etc. Currently, the wave-function is not symmetry enforced, thus translational invariance is not preserved during optimizations. For PBC, the energy precision can be further improved by enforcing translational invariance on the wave-function. The Tensor-Backflow has no restrictions on lattice shapes and boundary conditions, thus it is potentially an efficient and universal method for solving lattice fermion systems.
Figure 9: Comparisons of ground state patterns between two optimization strategies in Sec.III.1. The case is nn=0.875, UU=8 and $t^{\prime}$=0 on the $4\times 16$ lattie under OBC, with NN backflow in wave-functions. Left-sided figures denote the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, and right-sided figures denote the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. (a)(b) and (c)(d) are achieved by optimization strategies 1 and 2, respectively.
Figure 10: Comparisons of structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$ from (a)(b)the initial UHF and (c)(d)the converged wave-function. The case is nn=0.875, UU=8 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC, with NN backflow in the wave-function.
@@ V Acknowledgement
Figure 11: Comparisons of structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$ from (a)(b)the initial UHF, (c)(d)the converged wave-function. The case is nn=0.875, UU=10 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC, with NN backflow in the wave-function.
Figure 12: Comparisons of structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$ from (a)(b)the initial UHF, (c)(d)the converged wave-function. The case is nn=0.875, UU=12 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC, with NN backflow in the wave-function.
Appendix A Comparing ground state patterns of two optimization strategies
Fig.(9) compares ground state patterns between two optimization strategies in Sec.III.1, and results are from pp=0 wave-functions. The case is nn=0.875, UU=8 on the $4\times 16$ lattice under OBC. From the figure, both optimizations strategies achieve similar ground state patterns. Meanwhile, ground state patterns in Fig.(9) yield to those from state-of-the-art NQS results Hubbard_NN3 .
Figure 13: Energy convergence for nn=0.875 and $t^{\prime}$=0 on the $16\times 16$ lattice when the wave-function is transferred from UU=8, with NN backflow in wave-functions. More steps are required when the case has larger difference than UU=8 and PBC. Converged energies are -0.6787, -0.6272 and -0.6786 in (a)(b) and (c), respectively.
Appendix B Comparisons between initial UHFs and subsequent Tensor-Backflow wave-functions
We show that Tensor-Backflow achieves an essentially different state from the initial UHF by structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$. For nn=0.875 on the $16\times 16$ lattice under PBC, cases of UU=8,10,12 are depicted in Fig.(10), Fig.(11) and Fig.(12), respectively. Based on definitions in Sec.III.3, correlations in both $C(\mathbf{k})$ and $S(\mathbf{k})$ are classical, thus they are calculated through averages on MC samples. Here, each structure factor is obtained by 46080 MC samples, results of UHF and Tensor-Backflow are from pp=0 and pp=1 wave-functions, respectively.
In our calculations, each MC sample is flattened from two-dimensional to one-dimensional in the same way and each struture factor is calcualted by the same procedure from MC samples. Comparing (a)(b) to (c)(d) in Fig.(10), (11) and (12), properties of the converged wave-function is not simply inherited from the initial UHF. Furthermore, the Tensor-Backflow can override properties of the initial UHF. In each case, there is $U_{\mathrm{UHF}}=U$, and initial UHFs are not fully optimized in all three cases.
Appendix C Energy convergence of transferring from UU=8
Fig.(13) depicts energy convergences when transferring the converged wave-function for UU=8 and PBC to other cases. Because of the good initial state, required optimization step number is much less than that from a direct optimization. For cases denoted by (a), (b) and (c) in Fig.(13), energies achieved by direct optimizations are -0.6771, -0.6264 and -0.6771, respectively. Meanwhile, energies achieved by transferring from UU=8 are -0.6787, -0.6272 and -0.6786, respectively.
Appendix D Comparing ground state patterns for various $U_{\mathrm{UHF}}$ for $t^{\prime}$=-0.2
For the case of nn=0.875, UU=8, $t^{\prime}$=-0.2 on the $12\times 12$ lattice under PBC, comparisons of ground state patterns of $U_{\mathrm{UHF}}$=0, 4 and 8 are depicted in Fig.(14), with converged energies from Tab.(6). All optimization are performed without pinning-fields applied.
From the figure, different periods of stripes are achieved with different $U_{\mathrm{UHF}}$ Hubbard_NN4 . Therefore, $U_{\mathrm{UHF}}$ is essential to avoid local minima when optimizing Tensor-Backflow. Based on the density $n_{i}$, the lowest energy of -0.7420 is achieved by the width-4 stripe, yielding to the state-of-the-art NQS result Hubbard_NN3 .
Figure 14: Ground state patterns for nn=0.875, UU=8, $t^{\prime}$=-0.2 on the $12\times 12$ lattice under PBC for various $U_{\mathrm{UHF}}$, with NNN backflow in wave-functions. Left-sided figures denote the averate spin per site: $s_{i}^{z}$ and right-sided figures denote the averate density per per site: $n_{i}$. Results of $U_{\mathrm{UHF}}$=0, 4, 8 are depicted in subfigures (a)(b), (c)(d) and (e)(f), respectively. All results are from pp=1 wave-functions. No pinning-fields are applied during optimizations.